OpenStudy (anonymous):

Tiana would like to buy ice cream cones for her birthday party, but she wants cones with the largest volume. Which of the following brands of ice cream cones should she buy? A. Brand A B. Brand B C. Either - the cones have equal volumes

OpenStudy (anonymous):

you dont give enough values to figure out the answer

OpenStudy (anonymous):

do you have any numbers for each brand?

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

are you aware of the formula for the volume of a cone?

OpenStudy (anonymous):

i am not

OpenStudy (anonymous):

the basic formula should be Volume of cone= 1/3bxh

OpenStudy (anonymous):

so all you have to do is plug in your numbers

OpenStudy (anonymous):

so for the first cone

OpenStudy (anonymous):

you would put V= 1/3 times 5 times 16

OpenStudy (anonymous):

if you're lost let me know

OpenStudy (anonymous):

so i believe its both

OpenStudy (anonymous):

@lilcrunch the formula for the ice cream cone is not correct

OpenStudy (anonymous):

\[V= \frac{ \pi r^{2} h }{ 3 }\] Therefore, plug your known radius and height for each cone to find the answer. Lemme know if you need more hints.

OpenStudy (anonymous):

V of ice cream A = ( pi 5^2 (16) ) /3 = 400pi/3 = 133.33 pi V of ice cream B = ( pi 4^2 (20) ) /3 = 320 pi/3 = 106. 67 pi As in comparison, you can figure out which one would give her more volume of ice cream :)